### v1

# CKGConv: General Graph Convolution with Continuous Kernels

Abstract
The existing definitions of graph convolution, either from spatial or spectral perspectives, are inflexible and not unified. Defining a general convolution operator in the graph domain is challenging due to the lack of canonical coordinates, the presence of irregular structures, and the properties of graph symmetries. In this work, we propose a novel graph convolution framework by parameterizing the kernels as continuous functions of pseudo-coordinates derived via graph positional encoding. We name this Continuous Kernel Graph Convolution (CKGConv). Theoretically, we demonstrate that CKGConv is flexible and expressive. CKGConv encompasses many existing graph convolutions, and exhibits the same expressiveness as graph transformers in terms of distinguishing non-isomorphic graphs. Empirically, we show that CKGConv-based Networks outperform existing graph convolutional networks and perform comparably to the best graph transformers across a variety of graph datasets.
1 Introduction
Recent advances in applying Transformer architectures in computer vision ignited a heated competition with the predominant Convolutional Neural Networks (ConvNets) He et al. (2016a); Tan and Le (2019). This rivalry started when Vision Transformers (ViTs) Dosovitskiy et al. (2020); Wang et al. (2021); Liu et al. (2021, 2022a) exhibited impressive empirical gains over the state-of-the-art ConvNet architectures of the time. However, several recent ConvNet variants Liu et al. (2022b); Woo et al. (2023) achieve similar performance as ViTs by incorporating innovative designs such as larger kernels and depthwise convolutions Chollet (2017).
In contrast, the appeal of Convolutional Graph Neural Networks (GNNs) seems to be diminishing; Graph Transformers (GTs) demonstrate elevated efficacy on many challenging graph learning tasks Ying et al. (2021); Rampášek et al. (2022); Zhang et al. (2023); Ma et al. (2023). One reason might be that most existing definitions of graph convolution are inflexible and not unified. Message-passing Neural Networks (MPNNs) Gilmer et al. (2017); Veličković (2022) are defined in the spatial domain and limited to a 1-hop neighborhood; Spectral GNNs Bruna et al. (2014) are defined from a graph-frequency perspective and require careful designs (e.g., polynomial approximation Defferrard et al. (2016); Wang and Zhang (2022) or transformer-encoder Bo et al. (2023)) to generalize to unseen graphs. Defining a general convolution operator in the graph domain is challenging due to the characteristics of a graph: the lack of canonical coordinates, the presence of irregular structures, and the graph symmetries. These aspects are fundamentally different from Euclidean spaces.
By addressing the aforementioned challenges, we generalize the continuous kernel convolution Romero et al. (2022) to graph domains, and propose a general convolution framework, namely Continuous Kernel Graph Convolution (CKGConv, as shown in Fig. 1). This subsumes several common convolutional GNNs, including non-dynamic [^1] We use the terms ‘dynamic’, ‘data-dependent’ and ‘feature-dependent’ interchangeably Park and Kim (2021). MPNNs, polynomial spectral GNNs, and diffusion-enhanced GNNs. We propose three designs to address the challenges of the graph domain: (1) pseudo-coordinates for graphs via positional encoding (addressing the lack of canonical coordinates and handling symmetries); (2) support-size invariant convolution (compensating for irregular structure); (3) adaptive degree-scalers (mitigating the impact of symmetries).
Figure 1: Continuous Kernel Graph Convolution (CKGConv)
CKGConv is endowed with several desirable properties. It exhibits immunity against over-smoothing and over-squashing. It encompasses Equivariant Set Neural Networks Segol and Lipman (2020) as a special case for the setting where there is no observed graph. Furthermore, we theoretically prove that CKGConv with generalized distance (GD) can be as powerful as Generalized Distance Weisfeiler-Lehman (GD-WL) Zhang et al. (2023) in graph isomorphism tests (and is thus more powerful than 1-WL). Our experiments demonstrate the effectiveness of the proposed framework. CKGConv ranks among the best performing models in a variety of graph learning benchmarks and significantly outperforms existing convolutional GNNs.
Our contributions are summarized as follows:
  • We propose a novel graph convolution framework based on pseudo-coordinates of graphs and continuous convolution kernels.
  • We demonstrate theoretically and empirically that our proposed graph convolution is expressively powerful, outperforming existing convolutional GNNs and achieving comparable performance to the state-of-the-art graph transformers.
  • Various exploratory studies illustrate that the capacity to generate both positive and negative kernel coefficients endows CKGConv with advantages from the graph signal processing perspective, and differentiates its behavior from graph transformers employing attention mechanisms. This brings up the potential direction to combine CKGConv and attention mechanisms, toward the more powerful graph models.
2 Related Work
Message-Passing Neural Networks MPNNs Gilmer et al. (2017); Veličković (2022) are a class of GNNs widely used in graph learning tasks. To update the representation of a node $i$, MPNNs aggregate the features from the neighbors of $i$. In most cases, the message-passing mechanisms can be viewed as convolution with kernels locally supported on a $1$-hop neighborhood. The kernels may be fixed Kipf and Welling (2017); Xu et al. (2019); Hamilton et al. (2017), learnable Monti et al. (2017), or dynamic (i.e., feature-dependent) Veličković et al. (2018); Corso et al. (2020); Hamilton et al. (2017). Recent Graph Rewiring techniques extend MPNNs beyond $1$-hop by introducing additional edges, guided by curvature Topping et al. (2022), spectral gap Arnaiz-Rodríguez et al. (2022), geodesic distance Gutteridge et al. (2023), or positional encoding (PE) Gabrielsson et al. (2023) . The usage of PE by Gabrielsson et al. (2023) has some similarity to our work, but is constrained by the MPNN framework and lacks flexibility.
Spectral Graph Neural Networks In contrast to MPNNs, Spectral Graph Neural Networks define graph filters [^2] We use ‘filter’ and ‘kernel’ interchangeably in this work, following the related literature on graph convolution. in the spectral domain. The pioneering approach by Bruna et al. (2014) cannot generalize to an unseen graph with a different number of nodes. Follow-up works, constituting the class of Polynomial Spectral Graph Neural Networks, address this by (approximately) parameterizing the spectral filters by a polynomial of the (normalized) graph Laplacian Defferrard et al. (2016); He et al. (2021); Liao et al. (2022); Wang and Zhang (2022). Similarly, Diffusion Enhanced Graph Neural Networks extend spatial filters beyond $1$-hop by a polynomial of diffusion operators (e.g., adjacency matrix, random walk matrix)  Gasteiger et al. (2019b, a); Chien et al. (2021); Zhao et al. (2021a); Frasca et al. (2020), exhibiting strong connections to spectral GNNs. Notably, besides the polynomial approach, recent works endeavor to generalize spectral GNNs by introducing graph-order invariant operations Beani et al. (2021); Bo et al. (2023).
Graph Transformers GTs are equipped with the transformer architecture Vaswani et al. (2017), consisting of self-attention mechanisms (SAs) and feed-forward networks. Due to the characteristics of the graph domain, directly migrating the transformer architecture to graph learning tasks leads to poor performance Dwivedi and Bresson (2021) due to the lack of topological information. Modern graph transformer architectures address this by integrating message-passing mechanisms Kreuzer et al. (2021); Chen et al. (2022); Rampášek et al. (2022) or incorporating graph positional encoding (PE) Kreuzer et al. (2021); Ying et al. (2021); Mialon et al. (2021); Hussain et al. (2022); Park et al. (2022); Kim et al. (2022); Zhang et al. (2023); Ma et al. (2023); Zhou et al. (2023). Appendix C.1 provides more detail about graph PE.
Expressiveness of Graph Neural Networks Graph isomorphism tests have been widely used to measure the theoretical expressiveness of GNNs in terms of their ability to encode topological patterns Xu et al. (2019). Without additional elements, MPNNs’ expressive power is bounded by first-order Weisfeiler-Lehman ($1$-WL) algorithms Xu et al. (2019). Polynomial spectral GNNs are as powerful as $1$-WL algorithms; it is not known if this is a bound. Higher-order GNNs Morris et al. (2019) are as powerful as $K$-WL 3332222-WL is equivalent to $1$-WL but have $O(N^{K})$ computational complexity. Recently, Zhang et al. (2023) demonstrate that Graph Transformers with generalized distance (GD) can go beyond $1$-WL but are still bounded by $3$-WL, with $O(N^{2})$ complexity.
Continuous Kernel Convolution in Euclidean Spaces Recently, to handle irregularly sampled data and data at different resolutions, Romero et al. (2022) and Knigge et al. (2023) propose learning a convolution kernel as a continuous function, parameterized by a simple neural network of the coordinates (relative positions), resulting in Continuous Kernel Convolution (CKGConv). This enables the convolution to generalize to any arbitrarily support size with the same number of learnable parameters. Driven by different motivations, several works have explored similar ideas for point cloud data Hermosilla et al. (2018); Wu et al. (2019); Xu et al. (2018); Li et al. (2018b); Hua et al. (2018). From a broader perspective, CKConv can be viewed as a subdomain of Implicit Neural Representation Mildenhall et al. (2020); Sitzmann et al. (2020); Tancik et al. (2020), where the representation targets are the convolution kernels. Note that these techniques rely on canonical coordinates and cannot be directly applied to non-Euclidean domains like graphs.
3 Methodology
3.1 Preliminary: Continuous Convolution Kernels
Let $x:\mathbb{Z}\to\mathbb{R}$ and $\psi:\mathbb{Z}\to\mathbb{R}$ be two scalar-valued real sequences sampled on the set of integers $\mathbb{Z}$, where $x[k]$ and $\psi[k]$ denote the signal and filter impulse response (kernel) at time $k$, respectively. The discrete convolution between the signal and the kernel at time $k$ is defined as follows:
$(x\star\psi)[k]:=\sum_{\ell\in\mathbb{Z}}x[\ell]\psi[k-\ell],$ (1)
In most cases, the kernel is of finite width $n_{\psi}$, i.e., $\psi[k]=0$ if $k<0$ or $k\geqslant n_{\psi}$. The convolution sum in Eq. (1) is accordingly truncated at $\ell\notin[k-n_{\psi}+1,k]$. However, learning such fixed support, discrete kernels cannot be generalized to arbitrary widths (i.e., different $n_{\psi}$) with the same set of parameters.
To address this shortcoming,  Romero et al. (2022) propose to learn convolutional kernels by parameterizing $\psi[k]$ via a continuous function of $k$, implemented using a small neural network (e.g., multi-layer perception (MLP)). This is termed Continuous Kernel Convolution (CKConv). This formulation allows CKConv to model long-range dependencies and handle irregularly sampled data in Euclidean spaces.
3.2 Graph Convolution with Continuous Kernels
In the graph domain, convolution operators are required to handle varying sizes of supports, due to a varying number of nodes and irregular structures. We explore the potential of continuous kernels for graphs and propose a general graph convolution framework with continuous kernels.
The generalization of continuous kernels to graph domain is not trivial due to the following characteristics: (1) the lack of canonical coordinates makes it difficult to define the relative positions between non-adjacent nodes; (2) the irregular structure requires the kernel to generalize to different support sizes while retaining numerical stability; (3) the presence of graph symmetries demands that the kernel can distinguish between nodes in the support without introducing permutation-invariant operations or ambiguities.
These challenges drive us to propose a general graph convolution framework with continuous kernels, namely CKGConv. Our overall design consists of three innovations: (1) we use graph positional encoding to derive the pseudo-coordinates of nodes and define relative positions; (2) we introduce a support-size invariant convolution to handle the irregular structure; (3) we incorporate an adaptive degree-scaler to improve the representation of structural graph information.
3.2.1 Graph Positional Encoding as Pseudo-coordinates
In contrast to Euclidean spaces, the graph domain is known to lack canonical coordinates. Consequently, it is not trivial to define the relative distance between two non-adjacent nodes. Pioneering work Monti et al. (2017) attempted to define pseudo-coordinates for graphs, but it was restricted to the adjacent nodes.
In this work, we reveal that graph positional encoding (PE) can be naturally employed to construct pseudo-coordinates, allowing us to specify relative positions for continuous kernels. Specifically, we use Relative Random Walk Probabilities (RRWP) Ma et al. (2023), which have been demonstrated to be one of the most expressive graph positional encodings. Let $\mathbf{A}\in\mathbb{R}^{N\times N}$ be the adjacency matrix of a graph $G=(\mathcal{V},\mathcal{E})$ with $N$ nodes, and let $\mathbf{D}$ be the diagonal degree matrix, $\mathbf{D}_{i,i}=\sum_{j\in\mathcal{V}}\mathbf{A}_{i,j}$. The random walk matrix is $\mathbf{M}:=\mathbf{D}^{-1}\mathbf{A}$. Entry $\mathbf{M}_{ij}$ is then the probability of a move from node $i$ to node $j$ in one step of a simple random walk. With $\mathbf{I}\in\mathbb{R}^{N\times N}$ denoting the identity matrix, the RRWP for each pair of nodes $i,j\in\mathcal{V}$ is defined as:
$\mathbf{P}_{i,j}=[\mathbf{I},\mathbf{M},\mathbf{M}^{2},\dots,\mathbf{M}^{K-1}]_{i,j}\in\mathbb{R}^{K},$ (2)
RRWP is not the only choice for constructing pseudo-coordinates. One can use other graph positional encodings such as shortest-path-distance (SPD) Ying et al. (2021) and resistance distance (RD) Zhang et al. (2023).
3.2.2 Support-size Invariant Convolution
Contrary to existing literature focusing on tasks in the Euclidean space, the irregular structure of graphs demands that the graph convolution is invariant to the size of the support. When applying a kernel to different nodes, the support size can differ, and it is important that this does not induce disproportionate scaling of different node representations.
Therefore, we equip the proposed CKGConv with a support-size invariance. We consider a kernel function $\psi:\mathbb{R}^{r}\to\mathbb{R}$. For a graph $G=(\mathcal{V},\mathcal{E})$ with node-signal function $\chi:\mathcal{V}\to\mathbb{R}$, CKGConv is defined as:
$(\chi\star\psi)(i):=\frac{1}{|\operatorname{supp}_{\psi}(i)|}\sum_{j\in\operatorname{supp}_{\psi}(i)}\chi(j)\cdot\psi(\mathbf{P}_{i,j})+b\,.$ (3)
Here $b\in\mathbb{R}$ is a learnable bias term; the set $\operatorname{supp}_{\psi}(i)$ is the predefined support of kernel $\psi$ for node $i$ (i.e., $|\operatorname{supp}_{\psi}(i)|$ denotes the kernel size); and $\mathbf{P}_{i,j}\in\mathbb{R}^{K}$ is the relative positional encoding.
We can set $\operatorname{supp}_{\psi}(i)$ to be the $K$-hop neighborhood  [^4] The set of nodes whose shortest-path distance to node $i$ is smaller and equal to $K$ of node $i$. Alternatively, we can choose the support to be the entire graph for a global kernel. This flexibility arises because the construction of pseudo-coordinates is decoupled from the evaluation of the convolution kernel. We show that the globally supported variant is endowed with several desired theoretical properties ( Sec. 3.3 and Sec. 4.2).
3.2.3 Depthwise Separable Convolution
We extend the scalar-valued definition of CKGConv (shown in Eq. (3)) to vector-valued signals ($\boldsymbol{\chi}:\mathcal{V}\to\mathbb{R}^{d}$ and $(\boldsymbol{\chi}\star\boldsymbol{\psi})(i):\mathcal{V}\to\mathbb{R}^{d^{\prime}}$) via the Depthwise Separable Convolution (DW-Conv) architecture Chollet (2017),
$(\boldsymbol{\chi}\star\boldsymbol{\psi})(i):=\mathbf{W}\big{(}\tfrac{1}{|\operatorname{supp}_{\psi}(i)|}\sum_{j\in\operatorname{supp}_{\psi}(i)}\boldsymbol{\chi}(j)\odot\boldsymbol{\psi}(\mathbf{P}_{i,j})\big{)}+\mathbf{b}\,.$ (4)
Here $\boldsymbol{\psi}:\mathbb{R}^{K}\to\mathbb{R}^{d}$ is a kernel function acting on a vector; $\mathbf{W}\in\mathbb{R}^{d^{\prime}\times d}$ and $\mathbf{b}\in\mathbb{R}^{d^{\prime}}$ are the learnable weights and bias, respectively, shared by all nodes; and $\odot$ stands for elementwise multiplication.
We can alternatively extend Eq. (3) to multiple channels via grouped convolution Krizhevsky et al. (2012), multi-head architectures Vaswani et al. (2017), or even MLP-Mixer architectures Tolstikhin et al. (2021); Touvron et al. (2023). We select DWConv because it provides a favorable trade-off between expressiveness and the number of parameters.
3.2.4 MLP-based Kernel Function
In this work, as an example, we introduce kernel functions parameterized by multi-layer perceptrons (MLPs), but the proposed convolution methodology accommodates many other kernel functions. Each MLP block consists of fully connected layers ($\mathtt{FC}$), non-linear activation ($\sigma$), a normalization layer ($\mathtt{norm}$) and a residual connection, inspired by ResNetv2 He et al. (2016b):
$\mathtt{MLP}(\mathbf{x}):=\mathbf{x}+\mathtt{FC}\circ\sigma\circ\mathtt{Norm}\circ\mathtt{FC}\circ\sigma\circ\mathtt{Norm}(\mathbf{x})\,.$ (5)
Here $\circ$ stands for the function composition; $\mathtt{FC}(\mathbf{x}):=\mathbf{W}\mathbf{x}+\mathbf{b}$, with learnable weight $\mathbf{W}\in\mathbb{R}^{r\times r}$ and bias $\mathbf{b}\in\mathbb{R}^{r}$, and we use GELU Hendrycks and Gimpel (2016) as the default choice of $\sigma$.
The overall kernel function $\boldsymbol{\psi}$ is defined as:
$\boldsymbol{\psi}(\mathbf{\mathbf{P}_{i,j}}):=\mathtt{FC}\circ\mathtt{Norm}\circ\mathtt{MLP}\circ\dots\circ\mathtt{MLP}(\mathbf{P}_{i,j})\,,$ (6)
where the last $\mathtt{FC}:\mathbb{R}^{r}\to\mathbb{R}^{d}$ maps to the desired number of output channels.
3.2.5 Degree Scaler
As a known issue in graph learning, average-based convolutions/aggregations cannot preserve the degree information of nodes Xu et al. (2019); Corso et al. (2020). Therefore, we introduce a post-convolution adaptive degree-scaler into the node representation to recover such information, following the approach propsed by Ma et al. (2023):
$\mathbf{x}_{i}^{\prime}:=\mathbf{x}_{i}\odot\boldsymbol{\theta}_{1}+\left(d_{i}^{1/2}\cdot\mathbf{x}_{i}\odot\boldsymbol{\theta}_{2}\right)\in\mathbb{R}^{d}\,.$ (7)
Here $d_{i}\in\mathbb{R}$ is the degree of node $i$, and $\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}\in\mathbb{R}^{r}$ are learnable weight vectors.
As an alternative, we also introduce a variant that injects the degree information directly into the RRWP $\mathbf{P}_{i,j}\in\mathbb{R}^{r}$ before applying the kernel function $\psi$:
$\mathbf{P}_{i,j}:=\mathbf{P}_{i,j}\odot\boldsymbol{\theta}_{1}+\left((d_{i}^{1/2}\odot\boldsymbol{\theta}_{2})\odot\mathbf{P}_{i,j}\odot(d_{j}^{-1/2}\odot\boldsymbol{\theta}_{3})\right)\,,$ (8)
where $\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2},\boldsymbol{\theta}_{3}\in\mathbb{R}^{r}$ are learnable parameters. This variant enjoys several desired theoretical properties as discussed in Sec. 4.3. However, in practice, we did not observe any significant differences in empirical performance, and use Eq. (7) in our experiments due to its computational efficiency.
3.2.6 Overall Architecture of CKGCN
The overall multi-layer architecture of the proposed model, denoted by CKGCN, consists of $L$ CKGConv-blocks as the backbone, together with task-dependent output heads (as shown in Fig. 5 in Appendix A). Each CKGConv block consists of a CKGConv layer and a feed-forward network (FFN), with residual connections and a normalization layer:
$\mathtt{CKGConvBlock}(\cdot):=\mathtt{norm}\circ\mathtt{FFN}\circ\mathtt{norm}\circ\mathtt{CKGConv}(\cdot)\,.$ (9)
We use BatchNorm Ioffe and Szegedy (2015) in the main branch as well as in the kernel functions. Using LayerNorm Ba et al. (2016) has the potential to cancel out the degree information Ma et al. (2023) Appendix. D.4) presents additional architectural details. The input node/edge attributes ($\mathbf{x}^{\prime}_{i}\in\mathbb{R}^{d_{h}}$ and $\mathbf{e}^{\prime}_{i,j}\in\mathbb{R}^{d_{e}}$) and the absolute/relative positional encoding (RRWP) are concatenated: $\mathbf{x}_{i}=[\mathbf{x}^{\prime}_{i}\|\mathbf{P}_{i,i}]\in\mathbb{R}^{d_{h}+K}$ and $\mathbf{P}_{i,j}=[\mathbf{e}^{\prime}_{i,j}\|\mathbf{P}_{i,j}]\in\mathbb{R}^{d_{e}+K}$. A linear projection (a stem) maps to the desired dimensions before the backbone. If the data does not include node/edge attributes, zero-padding is used. For a fair comparison, we use the same task-dependent output heads as previous works Rampášek et al. (2022); Ma et al. (2023).
3.3 Theory: CKGConv Is as Powerful as Graph Transformers with Generalized Distance
Zhang et al. (2023) prove that Graph Transformers with generalized distance (GD) can be as powerful as GD-WL with a proper choice of attention mechanisms, thus going beyond 1-WL. We provide a similar result, demonstrating that CKGConv with GD is as powerful as GD-WL. It achieves the same theoretical expressiveness as SOTA graph transformers Zhang et al. (2023); Ma et al. (2023), with respect to the GD-WL test.
Proposition 3.1.
A CKGCN, stacking feed-forward networks (FFNs) and globally supported CKGConvs with generalized distance (GD) as pseudo-coordinates, is as powerful as GD-WL, when choosing the proper kernel $\boldsymbol{\psi}$.
The proof is provided in Appendix D.1.
4 Relationship with Previous Work
4.1 Beyond the Limitations of MPNNs
Despite being widely used, MPNNs are known to exhibit certain limitations: (1) over-smoothing; (2) over-squashing and under-reaching; (3) expressive power limited to $1$-WL. By contrast, CKGConv inherently addresses these constraints. Over-smoothing Li et al. (2018a); Oono and Suzuki (2020) arises because most MPNNs apply a smoothing operator (a blurring kernels or low-pass filter). CKGConv can generate sharpening kernels, as illustrated in our experiments. Over-squashing Alon and Yahav (2020); Topping et al. (2022) is mainly due to the local message-passing. CKGConv has kernels whose support can go beyond $1$-hop neighborhood (extending even to the entire graph). We have demonstrated in Sec. 3.3 that CKGConv can reach expressive power equivalent to GD-WL, thus going beyond $1$-WL algorithms.
4.2 Equivariant Set Neural Networks
In general, graphs can be viewed as a set of nodes with observed structures among them. Here, we demonstrate that when the pseudo-coordinates do not encode any graph structure, CKGConv can naturally degenerate to the general form of a layer in an Equivariant Set Network Segol and Lipman (2020). This matches the natural transition between graph data and set data. The following proposition states that wen we use a first-order RRWP to generate the positional encodings (and thus ignore any graph structure), CKGConv is equivalent to a layer of an equivariant set network.
Proposition 4.1.
With $1$-order RRWP $\mathbf{P}=[\mathbf{I}]$, CKGConv with a globally supported kernel can degenerate to the following general form of a layer in an Equivariant Set Network (Eq. (8) in  Segol and Lipman (2020)):
$(\chi\star\psi)(i)=$ $\gamma\cdot\chi(i)+\beta\cdot\big{(}\frac{1}{|\mathcal{V}|}\sum_{j\in\mathcal{V}}\chi(j)\big{)}+b\,.$ (10)
Here $\gamma,\beta,b\in\mathbb{R}$ are learnable parameters. This can be directly generalized to vector-valued signals.
The proof is provided in Appendix D.2.
4.3 Polynomial Spectral and Diffusion Enhanced GNNs
Polynomial spectral GNNs Defferrard et al. (2016); He et al. (2021); Wang and Zhang (2022) apply convolutional kernels expressed as fixed-order polynomial functions of the symmetric normalized graph Laplacian matrix $\tilde{\mathbf{L}}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}\in\mathbb{R}^{n\times n}$. Furthermore, diverse Diffusion Enhanced GNNs Gasteiger et al. (2019a); Frasca et al. (2020); He et al. (2021); Wang and Zhang (2022) use a similar polynomial parameterization with the diffusion operator $\tilde{\mathbf{A}}$ or $\mathbf{M}$ replacing the graph Laplacian.
The following proposition states that CKGConv can represent any polynomial spectral GNN or diffusion enhanced GNN of the same order.
Proposition 4.2.
With $K{-}1$-th order of RRWP $\mathbf{P}_{i,j}\in\mathbb{R}^{K}$ as pseudo-coordinates, CKGConv with a linear kernel $\psi$ can represent any Polynomial Spectral GNN or any Diffusion Enhanced GNNs of the same order exactly, regardless of the specific polynomial parameterization, if degree $d_{i}^{1/2}$ and $d_{j}^{-1/2}$ are injected to $\mathbf{P}_{i,j}$ properly.
The proof is presented in Appendix D.3. If $K\to\infty$, CKGConv can closely approximate a full spectral GNNs. This also highlights the relationship between RRWP and graph spectral wavelets  Hammond et al. (2011).
We note that CKGConv is strictly more expressive than polynomial spectral GNNs and diffusion-enhanced GNNs. Continuous convolution using any non-linear kernel $\psi$ cannot be represented by a polynomial spectral GNN or a diffusion-enhanced GNN. Since MLPs are universal function approximators Hornik et al. (1989), CKGConv can represent a considerably richer class of functions.
4.4 Fourier Features of Graphs
As mentioned in Proposition 4.2, RRWP can be viewed as a set of bases for a polynomial vector space, which approximates the full Fourier basis of graphs (i.e., eigenvectors of the Laplacian). Therefore, RRWP can be viewed as a set of (approximate) Fourier features under certain transformations. Likewise in the Euclidean space, Tancik et al. (2020) propose to construct Fourier features from coordinates to let MLPs better capture high-frequency information.
Specformer Bo et al. (2023) generates the graph spectral filters via transformers given the full or sampled Fourier bases in the spectral domain. This is generalizable to graphs with different orders. Specformer is the sibling work of CKGConv from the spectral perspective. In a similar fashion to the contrast between the Fourier transform and a wavelet decomposition, Specformer achieves better spectral localization whereas CKGConv exhibits better localization spatially.
4.5 Graph Transformers
As shown in Sec. 3.3, with the same generalized distances (e.g., SPD, RD, RRWP) as relative positional encoding or pseudo-coordinates, CKGConv can reach the same theoretical expressive power as Graph Transformers, with respect to graph isomorphism tests. However, good performance at graph isomorphism tests does not necessarily translate to expressiveness in real-world graph learning tasks.
From a filtering perspective, self-attention in (Graph) Transformers can be viewed as a feature-dependent filter Park and Kim (2021). However, the elements of the filter are constrained to be positive and thus the filter can only act as a blurring kernel, performing low-pass smoothing. In contrast, CKGConv is a feature-independent filter, but has the flexibility to include positive and negative coefficients. It can thus learn blurring, sharpening, or identity kernels.
In this work, we are not claiming that CKGConv is better than Graph Transformers or vice versa. We aim to emphasize that each approach has its own advantages. The contrasting strengths of feature-dependence and high-pass or sharpening, present an intriguing possibility of developing architectures that combine the strengths of graph transformers and continuous convolution.
5 Experimental Results
Table 1: Test performance in five benchmarks from Dwivedi et al. (2022a); Ma et al. (2023); Bo et al. (2023). Shown is the mean $\pm$ s.d. of 4 runs with different random seeds. Highlighted are the top first, second, and third results. # Param under $500K$ for ZINC, PATTERN, CLUSTER and $\sim 100K$ for MNIST and CIFAR10.
Model ZINC MNIST CIFAR10 PATTERN CLUSTER
[…]
GCN $0.367\pm 0.011$ $90.705\pm 0.218$ $55.710\pm 0.381$ $71.892\pm 0.334$ $68.498\pm 0.976$
GIN $0.526\pm 0.051$ $96.485\pm 0.252$ $55.255\pm 1.527$ $85.387\pm 0.136$ $64.716\pm 1.553$
GAT $0.384\pm 0.007$ $95.535\pm 0.205$ $64.223\pm 0.455$ $78.271\pm 0.186$ $70.587\pm 0.447$
GatedGCN $0.282\pm 0.015$ $97.340\pm 0.143$ $67.312\pm 0.311$ $85.568\pm 0.088$ $73.840\pm 0.326$
GatedGCN-LSPE $0.090\pm 0.001$- -- -- --
PNA $0.188\pm 0.004$ $97.94\pm 0.12$ $70.35\pm 0.63$- --
GSN $0.101\pm 0.010$- -- -- --
DGN $0.168\pm 0.003$- ${72.838\pm 0.417}$ $86.680\pm 0.034$-
Specformer $\mathbf{0.066\pm 0.003}$
CIN ${0.079}\pm{0.006}$- -- -- --
CRaW1 $0.085\pm 0.004$ ${97.944}\pm{0.050}$ $69.013\pm 0.259$- --
GIN-AK+ ${0.080}\pm{0.001}$- $72.19\pm 0.13$ $\mathbf{86.850\pm 0.057}$-
SAN $0.139\pm 0.006$- -- $86.581\pm 0.037$ $76.691\pm 0.65$
Graphormer $0.122\pm 0.006$- -- -- --
K-Subgraph SAT $0.094\pm 0.008$- -- ${86.848\pm 0.037}$ $77.856\pm 0.104$
EGT $0.108\pm 0.009$ $\mathbf{98.173\pm 0.087}$ $68.702\pm 0.409$ $86.821\pm 0.020$ $\mathbf{79.232\pm 0.348}$
Graphormer-URPE $0.086\pm 0.007$- -- -- --
Graphormer-GD $0.081\pm 0.009$- -- -- --
GPS $\mathbf{0.070}\pm\mathbf{0.004}$ ${98.051\pm 0.126}$ ${72.298\pm 0.356}$ $86.685\pm 0.059$ ${78.016\pm 0.180}$
GRIT $\mathbf{0.059\pm 0.002}$ $\mathbf{98.108\pm 0.111}$ $\mathbf{76.468\pm 0.881}$ $\mathbf{87.196\pm 0.076}$ $\mathbf{80.026\pm 0.277}$
CKGCN $\mathbf{0.059\pm 0.003}$ $\mathbf{98.423\pm 0.155}$ $\mathbf{72.785\pm 0.436}$ $\mathbf{88.661\pm 0.143}$ $\mathbf{79.003\pm 0.140}$
5.1 Benchmarking CKGCN
We evaluate our proposed method on five datasets from Benchmarking GNNs Dwivedi et al. (2022a) and another two datasets from Long-Range Graph Benchmark Dwivedi et al. (2022b). These benchmarks include diverse node and graph level learning tasks such as node classification, graph classification, and graph regression. They test an algorithm’s ability to focus on graph structure encoding, to perform node clustering, and to learn long-range dependencies. The statistics of these datasets and further details of the experimental setup are deferred to Appendix B.
Baselines We compare our methods with the SOTA Graph Transformer: GRIT Ma et al. (2023); the hybrid Graph Transformer (MPNN+self-attention): GraphGPS Rampášek et al. (2022); popular Message-passing Neural Networks: GCN Kipf and Welling (2017), GIN  Xu et al. (2019) and its variant with edge-features Hu et al. (2020), GAT Veličković et al. (2018), GatedGCN Bresson and Laurent (2017), GatedGCN-LSPE Dwivedi et al. (2021), and PNA Corso et al. (2020); Other Graph Transformers: Graphormer Ying et al. (2021), K-Subgraph SAT Chen et al. (2022), EGT Hussain et al. (2022), SAN Kreuzer et al. (2021), Graphormer-URPE Luo et al. (2022), and Graphormer-GD Zhang et al. (2023); SOTA Spectral Graph Neural Networks: (Specformer Bo et al. (2023) and DGN Beani et al. (2021)); and other SOTA Graph Neural Networks: GSN Bouritsas et al. (2022), CIN Bodnar et al. (2021), CRaW1 Toenshoff et al. (2021), and GIN-AK+ Zhao et al. (2021b).
Benchmarks from Benchmarking GNNs In Table 1, we report the results on five datasets from Dwivedi et al. (2022a): ZINC, MNIST, CIFAR10, PATTERN, and CLUSTER. We observe that the proposed CKGConv achieves the best performance for 3 out of 5 datasets and is ranked within the three top-performing models for the other 2 datasets. Compared to the hybrid transformer, GraphGPS, consisting of an MPNN and self-attention modules (SAs), CKGCN outperforms on all five datasets, indicating the advantage of the continuous convolution employed by CKGConv over MPNNs, even when enhanced by self-attention. GRIT and CKGConv achieve comparable performance but exhibit advantage in different datasets. CKGConv outperforms on MNIST and PATTERN while GRIT performs better on CIFAR10 and CLUSTER. This suggests that the capability to learn dynamic kernels and sharpening kernels might have different impact and value on the empirical performance depending on the nature of the dataset. Notably, CKGConv exhibits superior performance compared to all other convolution-based GNNs for four of the five datasets. The only exception is CIFAR10, where CKGConv is slightly worse than DGN, although the difference is not statistically significant 555according to a two-sample one-tailed t-test..
Long-Range Graph Benchmark (LRGB) Graph transformers demonstrate advantages over MPNNs in modeling long-range dependencies. Here, we verify the capacity of CKGConv to model long-range dependencies. We conduct experiments on two peptide graph datasets from the Long-Range Graph Benchmark (LRGB) Dwivedi et al. (2022b). The obtained results are summarized in Table 2.
On both datasets, CKGConv obtains the second-best mean performance. Based on a two-sample one-tailed t-test, the performance is not significantly different from the best performing algorithm (GRIT). There is however, a statistically significant difference between CKGConv’s performance and the third best algorithm’s performance for both datasets. This demonstrates that our model is able to learn long-range interactions, on par with the SOTA graph transformers.
Table 2: Test performance on two benchmarks from long-range graph benchmarks (LRGB) Dwivedi et al. (2022b). Shown is the mean $\pm$ s.d. of 4 runs with different random seeds. Highlighted are the top first, second, and third results. # Param $\sim 500K$.
Model Peptides-func Peptides-struct
[…]
GCN $0.5930\pm 0.0023$ $0.3496\pm 0.0013$
GINE $0.5498\pm 0.0079$ $0.3547\pm 0.0045$
GatedGCN $0.5864\pm 0.0035$ $0.3420\pm 0.0013$
GatedGCN+RWSE $0.6069\pm 0.0035$ $0.3357\pm 0.0006$
Transformer+LapPE $0.6326\pm 0.0126$ $0.2529\pm 0.0016$
SAN+LapPE $0.6384\pm 0.0121$ $0.2683\pm 0.0043$
SAN+RWSE ${0.6439\pm 0.0075}$ $0.2545\pm 0.0012$
GPS $\mathbf{0.6535\pm 0.0041}$ $\mathbf{0.2500\pm 0.0012}$
GRIT $\mathbf{0.6988\pm 0.0082}$ $\mathbf{0.2460\pm 0.0012}$
CKGCN $\mathbf{0.6952\pm 0.0068}$ $\mathbf{0.2477\pm 0.0018}$
5.2 Advantages of Flexible Kernels with Positive and Negative Coefficients
Convolutional kernels with both negative and positive coefficients have a long history. In contrast to blurring/denoising kernels (all positive) to smooth the signals, such kernels are widely used to amplify the signal differences among data points, e.g., for image sharpening and ridge detection in image processing.
In this section, we illustrate that such kernels are also meaningful and advantageous in graph learning, with two toy examples on Anti-oversmoothing and Ridge-detection.
5.2.1 Toy Example: CKGConv Can Mitigate Oversmoothing
With the ability to generate both positive and negative coefficients, CKGConv can learn sharpening kernels (a.k.a. high-pass filters), which amplifies the signal differences among data points for alleviating oversmoothing. Here, we provide a toy example to better illustrate CKGConv’s capability to prevent oversmoothing.
We consider a simple graph with node signals as shown in Fig. 3, and train 2-layer and 6-layer GCNs and CKGCNs, with a 5-order RRWP, to predict labels that match the node signals. In this toy example, we remove all normalization layers, dropouts, as well as residual connections. All models are trained for 200 epochs with the Adam optimizer (initial learning rate 1e-3) to overfit this binary classification task. We report the results of 5 trials with different random seeds in Table 3.
Table 3: Anti-oversmoothing Toy Example (Fig. 3): Train performance on node signals reconstruction. Shown in the mean $\pm$ s.d. of 5 runs with different random seeds.
Train 2-Layer GCN 6-Layer GCN 2-Layer CKGCN 6-Layer CKGCN
BCE Loss 0.071 $\pm$ 0.044 0.693 $\pm$ 2e-05 4e-05$\pm$ 2e-05 0.0 $\pm$ 0.0
Accuracy (%) 100 $\pm$ 0 50 $\pm$ 0 100 $\pm$ 0 100 $\pm$ 0
As shown in the results, both the 2-layer GCN and 2-layer CKGCN can overfit the toy example and reach 100% accuracy. However, a 6-layer GCN fails to reconstruct the node signals. Applying 6 smoothing convolutions (all-positive filter coefficients) in this small network leads to the aggregated representation at each node being very similar. This is a typical oversmoothing effect. The network predicts all nodes to have the same label, resulting in 50% accuracy in the toy example. In contrast, a 6-layer CKGCN not only reaches 100% accuracy but also achieves a lower BCE loss, showcasing its strong capability in mitigating oversmoothing.
Node Signals 001111001111001111
Figure 2: The toy example for Anti-oversmoothing.
Node Signals 111111110000111111110000
Node Labels 001111111100001111111100
Figure 3: The toy example for Ridge Detection.
5.2.2 Toy Example: CKGConv Can Do Ridge Detection
In the graph domain, kernels with positive and negative coefficients can be used to detect the nodes with cross-community-connection (the border nodes). In such a setting, it is essential that the sign of the coefficient for the central node is opposite to the signs of the first-hop neighbors, in order to detect differences in attributes.
We introduce a toy example to demonstrate it as shown in Fig. 3: given a graph with simple scalar node signals that match the “community” of a node (0 or 1), the goal is to identify the border nodes with node labels as 1,
In this study, we consider single-channel convolution kernels. We compare CKGConv with three all-positive kernels: GCNConv Kipf and Welling (2017), CKGConv+Softmax (attention-like), and CKGConv+Softplus. CKGConv and its variants use 5-order RRWP with the hidden dimension as 5 in the kernel function. CKGCN+Softmax (sum-aggregation) and CKGCN+Softplus (mean-aggregation) apply Softmax and Softplus on the kernel coefficients, respectively, to constrain the kernels to have positive coefficients only.
We aim to verify the upper bounds for the expressivity of the convolutions by training them to overfit the task. Each convolution operator is trained for 200 epochs with the Adam optimizer (learning rate 1e-2) using binary cross entropy loss (BCE loss). We report the last training BCE loss and accuracy from 5 trials with different initialization, in Table 4.
Table 4: Ridge-detection Toy Example (Fig. 3): Train performance on node signals reconstruction. Shown in the mean $\pm$ s.d. of 5 runs with different random seeds.
Train CKGConv GCNConv CKGConv+Softmax CKGConv+Softplus
BCE Loss 2e-4 $\pm$ 1e-05 0.693 $\pm$ 0.001 0.693 $\pm$ 0 0.687 $\pm$ 0.049
Accuracy (%) 100 $\pm$ 0 50 $\pm$ 0 50 $\pm$ 0 60 $\pm$ 12.25
From the results, it is obvious that only convolution kernels with negative and positive values (regular CKConv) can reach 100% accuracy and achieve a low BCE loss. The other convolution kernels, with only positive values, all fail to identify the border nodes.
This toy example explains why negative coefficients are advantageous in some graph learning tasks.
5.2.3 CKGConv Kernel Visualization in Real-world Graph Learning
We introduce the flexibility of CKGConv to learn negative kernel coefficients from the data, but do not force CKGConv to learn negative kernel coefficients. Therefore, we visualize the learned kernels of CKGConv to showcase that CKGConv indeed learns to generate sharpening kernels when processing real-world graph datasets.
Specifically, we visualize two selected learned kernels from depthwise convolution of CKGConv for two graphs from the ZINC datasets, as shown in Fig. 4. We observe that in contrast to attention scores, forced to be all positive, in graph transformers, several learned kernels in CKGConv indeed include both positive and negative values, exhibiting the necessity of the flexible kernels.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: Adjacency matrices and learned continuous kernels across multiple channels for two graphs from the ZINC dataset
5.3 Sensitivity and Ablation Study of CKGConv
5.3.1 Sensitive study on the choice of graph PEs
In this paper, in order to avoid PE becoming the bottleneck of the model performance, we adopt one of the most expressive PEs, RRWP, from Ma et al. (2023). However, CKGConv is not constrained to working with a specific graph PE. As the PE summarizes the structural/positional information obtained from the graph topology in our CKGConv framework, the choices of PE affect the expressive power of CKGConv, as depicted in Proposition D.1 in our paper. Therefore, in this section, we study the impact of using different graph PEs, and showcase that CKConv can reach a competitive performance with other well-designed and expressive PEs besides RRWP.
We conduct the sensitivity study on ZINC datasets with four typical graph PEs: RRWP Ma et al. (2023), Resistance Distance (RD) Zhang et al. (2023), Shortest-path distance (SPD) Ying et al. (2021), and “Pair-RWSE", which is constructed as relative PE by concatenating the (absolute) Random Walk Structural Encoding (RWSE) Dwivedi et al. (2021) for each node-pair. Since RRWP inherently injects RWSE to node attributes, for a fair comparison, we add RWSE as the absolute PE to the node attribute while using other PEs. The experimental setup follows the main experiment and the results of 4 runs are reported in Table  5.
Table 5: The sensitive study of CKGCN on the choices of graph PEs. Shown is the mean ± s.d. of 4 runs with different random seeds.
CKGCN + RRWP RD SPD Pair-RWSE
MAE $\downarrow$ 0.059 $\pm$ 0.003 0.062 $\pm$ 0.004 0.072 $\pm$ 0.003 0.081 $\pm$ 0.002
The results of SPD and Pair-RWSE show that a sub-optimal PE design leads to worse performance of CKGCN. However, with an expressive PE, CKGCN demonstrates stable performance: CKGCN with both RD and RRWP reach competitive performance with state-of-the-art, which are not significantly different at the 5% level according to a two-sided t-test.
5.3.2 Sensitivity study on kernel support sizes of CKGConv
CKGConv framework allows the kernels with the pre-determined non-global supports, analogous to the regular convolution in Euclidean spaces. In this section, we study the effect of the different pre-determined support sizes based on $K$-hop neighborhoods on ZINC datasets. The sensitivity study follows the same experimental setup as the main experiments. The timing is conducted on a single a single NVIDIA V100 GPU (Cuda 11.8) and 20 threads of Intel(R) Xeon(R) Gold 6140 CPU @ 2.30GHz.
Table 6: The sensitivity study of the support sizes ($K$-hop neighborhoods) of CKGConv kernels with $21$-order RRWP. ($S$: kernel sizes; $N$: graph orders.)
ZINC MAE $\downarrow$ Run-Time (sec/ep) GPU-Mem (MB)
1-hop 0.073 $\pm$ 0.005 33.1 1186
3-hop 0.063 $\pm$ 0.002 33.6 1522
5-hop 0.061 $\pm$ 0.004 35.2 1624
11-hop 0.063 $\pm$ 0.002 33.2 2128
21-hop 0.060 $\pm$ 0.002 34.8 2148
Full 0.059 $\pm$ 0.003 35.4 2148
From the results, with the same order of RRWP, the larger support sizes usually lead to better empirical performances as well as the larger GPU memory consumption. On the one hand, the results showcase the performance stability of CKGConv, that all CKGCN variants with $K$-hop, $K>1$, reach competitive performance among the existing graph models. They outperform all existing GNNs and most Graph Transformers in the comparisons. On the other hand, the results justify the necessity of introducing graph convolutions beyond the 1-hop neighborhood (a.k.a., message-passing), given the fact that $1$-hop variant is significantly worse than the other variants with larger kernels. Furthermore, the sensitivity study also highlights the flexibility of CKGConv framework in balancing the computational cost and the capacity to model long-range dependencies, by effortlessly controlling the kernel sizes like the Euclidean convolutions. Note that the kernel sizes are not necessarily tied with the order of RRWP or the counterparts of other graph PEs in CKGConv.
5.3.3 Ablation study on the kernel functions.
As depicted in Sec. 4.3, polynomial-based GNNs can be viewed as CKGCN with linear kernel functions. However, allowing kernel functions with non-linearity is essential, since multilayer perceptions (MLPs) with non-linear activations can be universal function approximators Hornik et al. (1989) while linear functions cannot achieve that.
To better understand the effects of the choices of kernel functions, we conduct an ablation study on ZINC and PATTERN datasets, following the experimental setup of the main experiments. We compare CKGCN variants kernel functions with $0,1,2$ MLP-blocks (as shown in Eq. (5)). The width of each variant is adjusted to reach the parameter budget under $500$ K. Note that 00 MLP-block is equivalent to the linear kernel function and $2$ MLP-blocks setting is the default in CKGCN.
Table 7: The ablation study of # MLP blocks in the kernel function of CKGConv
# MLP ZINC PATTERN
Blocks MAE $\downarrow$ # param. W.Accuracy $\uparrow$ # param.
0 0.074 $\pm$ 0.005 487 K 87.355 $\pm$ 0.230 495 K
1 0.065 $\pm$ 0.005 438 K 88.955 $\pm$ 0.251 444 K
2 0.059 $\pm$ 0.003 434 K 88.661 $\pm$ 0.142 438 K
From the results of the ablation study (as shown in Table 7), CKGCNs with linear kernel functions under-perform the variants with non-linear kernel functions on both ZINC and PATTERN datasets, even with more learnable parameters. This observation matches our hypothesis on the indispensability of the non-linearity in kernel functions. It also justifies the advantage of CKGConv framework over the previous polynomial GNNs which can only introduce linear kernel functions.
5.4 CKGCN and GTs Behave Differently
In this section, we demonstrate that CKGCN shows significantly different behaviors compared to graph transformers in graph learning. This reveals the remarkable differences between GCKGConv and self-attention (SA), and indicates the potential advantages of proper combination of them in the model architecture.
5.4.1 Comparing CKGConv (Non-Dynamic-Sharpening) and SA (Dynamic-Blurring) kernels on Graphs
From a filtering perspective, SA in graph transformers can generate dynamic kernels but these are constrained to be blurring (low pass). By contrast, CKGConv kernels are non-dynamic but are capable of being sharpening (high-pass or band-pass).
In this experiment, we aim to understand whether these properties are both beneficial for graph learning tasks. Therefore, we introduce a variant of CKGConv that constrained to produce a blurring/low-pass kernel, namely CKGConv-Blur, by adding a Softmax on the kernels. This can be viewed as a blurring variant of CKGConv or a non-dynamic variant of self-attention. We conduct the comparison experiment on ZINC and PATTERN, on which CKGConv demonstrates comparable or better performance than GRIT.
Table 8: Effect of different kernel types on ZINC and PATTERN
Model ZINC PATTERN Dynamic Sharpened
[…]
GRIT $0.059\pm 0.002$ $87.196\pm 0.076$
CKGCN $0.059\pm 0.003$ $88.661\pm 0.143$
    -Blur $0.073\pm 0.003$ $87.000\pm 0.002$
From Table 8, we observe that from CKGConv-Blur, both adding either the dynamic property or the sharpening property can improve performance. In the case of ZINC, the performance improvement is dramatic. For PATTERN, dynamic kernels (GRIT) provide a small improvement, but sharpening kernels offer a more significant boost. PATTERN graphs are derived using stochastic block models, so we can conjecture that there is value in being able to construct a differentiating filter that is positive (smoothing) within the block of the node of interest and negative outside.
Note that SAs typically require incorporation of Softmax (or an approximation) to stabilize the attention scores. Thus, similar to recent work in the computer vision domain Park and Kim (2021), exploring how to combine CKGConvs and self-attention for graph models is a promising future research direction.
5.4.2 Ensemble of CKGCN and GRIT
In order to understand whether CKGConvs and self-attention learn complementary features, we conduct an ensembling experiment on ZINC to examine the effects of naively combining a SOTA graph transformer GRIT with CKGCN.
In Table 9, we report the MAE of an ensemble of GRIT models, an ensemble of CKGConv models, and a mixed ensemble using both of these models. In each case, the total ensemble size is 4 (two of each model for the mixed ensemble). Bootstrapping is employed to compute the mean and standard deviation of the MAE. We observe that constituting the ensemble using both CKGConv and SAs offers a statistically significant advantage compared to either homogeneous ensemble.
Table 9: Effect of ensembling on ZINC
Model GRIT-Ens. CKGConv-Ens. Mixed-Ens.
MAE $\downarrow$ 0.054$\pm$0.001 0.054$\pm$0.002 0.051$\pm$0.001*
We hypothesize that both CKGConvs and SAs have their own merits, and it can be further advantageous to suitably combine them in the model architecture. Similar efforts have been undertaken in computer vision Park and Kim (2021); Xiao et al. (2021).
6 Limitation
On the computational side, a naive implementation of CKGConv with global support requires $O(N^{2})$ complexity, the same as graph transformers. A more efficient implementation (Appendix B.7) can reduce the complexity to $O(NS)$ where $N:=|\mathcal{V}|$ and $S:=\mathbb{E}_{i\sim\mathcal{V}}[|\{\mathbf{P}_{i,j}\neq\mathbf{0}\}|]$. The complexity thus depends on the order of the RRWP and the graph structure. The implementation can prevent the usage of some operators such as BatchNorm.
7 Conclusion
Motivated by the lack of a flexible and powerful convolution framework for graphs, we propose a general graph convolution framework, CKGConv, by generalizing continuous kernels to graph domains. These can recover most non-dynamic convolutional GNNs, from spatial to (polynomial) spectral. Addressing the fundamentally different characteristics of graph domains from Euclidean domains, we propose three theoretically and empirically motivated designs to accomplish the generalization to graphs. Theoretically, we demonstrate that CKGConv possesses equivalent expressive power to SOTA Graph Transformers in terms of distinguishing non-isomorphic graphs via the GD-WL test Zhang et al. (2023). We provide theoretical connections to previous convolutional GNNs. Empirically, the proposed CKGConv architecture either surpasses or achieves performance comparable to the SOTA across a wide range of graph datasets. It outperforms all other convolutional GNNs and achieves performance comparable to SOTA Graph Transformers. A further exploratory experiment suggests that CKGConv can learn non-dynamic sharpening kernels and extracts information complementary to that learned by the self-attention modules of Graph Transformers. This motivates a potential novel avenue of combining CKGConv and SAs in a single architecture.
Appendix A Model Architecture and Implementation Details
A.1 Model Architecture
In order to put all the building blocks into one clear visualization, here we provide an overview of the CKGCN in Figure 5.
Figure 5: (a) Detailed Architecture of CKGCN with $L$ CKGConv blocks and task-dependent output head, (b) the detailed design of each CKGConv-block.
A.2 Preprocessing on Graph PEs
The expected values of random walk based graph PEs, e.g., RRWP, are dependent on the graph orders. For a graph with $N$ nodes, RRWP has the property that $\sum_{j\in\mathcal{V}}\mathbf{P}_{i,j}=\mathbf{1}\Rightarrow\mathbb{E}_{j\sim\mathcal{V}}[\mathbf{P}_{i,j}]=\mathbf{1}/N$. Empirically, we found that removing this dependency is beneficial to CKGConv. Therefore, we introduce an extra preprocessing for RRWP by setting $\mathbf{P}_{i,j}\leftarrow N\cdot\mathbf{P}_{i,j}$. For other graph PEs without such dependencies, e.g,. RD and SPD, this preprocessing is not necessary.
Appendix B Experimental Details
B.1 Description of Datasets
Table. 10 provides a summary of the statistics and characteristics of datasets used in this paper. The first five datasets are from Dwivedi et al. [2022a], and the last two are from Dwivedi et al. [2022b]. Readers are referred to Rampášek et al. [2022] for more details about the datasets.
Table 10: Overview of the graph learning datasets involved in this work Dwivedi et al. [2022a, b], Irwin et al. [2012].
Dataset # Graphs Avg. # nodes Avg. # edges Directed Prediction level Prediction task Metric
ZINC 12,000 23.2 24.9 No graph regression Mean Abs. Error
MNIST 70,000 70.6 564.5 Yes graph 10-class classif. Accuracy
CIFAR10 60,000 117.6 941.1 Yes graph 10-class classif. Accuracy
PATTERN 14,000 118.9 3,039.3 No inductive node binary classif. Weighted Accuracy
CLUSTER 12,000 117.2 2,150.9 No inductive node 6-class classif. Weighted Accuracy
Peptides-func 15,535 150.9 307.3 No graph 10-task classif. Avg. Precision
Peptides-struct 15,535 150.9 307.3 No graph 11-task regression Mean Abs. Error
B.2 Dataset splits and random seed
We conducted the experiments following the standard train/validation/test splits of the evaluated benchmarks, following previous works Rampášek et al. [2022], Ma et al. [2023]. For each dataset, we execute 4 runs with different random seeds (0,1,2,3) and report the mean performance and standard deviation.
B.3 Optimizer and Learning Rate Scheduler
We use AdamW Loshchilov and Hutter [2019] as the optimizer and the Cosine Annealing Learning Rate scheduler Loshchilov and Hutter [2017] with linear warm up as the learning rate scheduler.
B.4 Hyperparameters
Due to the limited time and computational resources, we did not perform an exhaustive search or a grid search on the hyperparameters. We mainly follow the hyperparameter setting of GRIT Ma et al. [2023], and make slight changes to adjust the number of parameters to match the commonly used parameter budgets.
We follow the most commonly used parameter budgets: up to 500k parameters for ZINC, PATTERN, CLUSTER, Peptides-func and Peptides-struct; and around 100k parameters for MNIST and CIFAR10.
The final hyperparameters are presented in Table 11 and Table 12.
Table 11: Hyperparameters for five datasets from BenchmarkingGNNs Dwivedi et al. [2022a]
Hyperparameter ZINC MNIST CIFAR10 PATTERN CLUSTER
# CKGConv-Block 10 4 3 10 16
   - Hidden dim 64 48 56 64 54
   - Dropout 0 0 0 0 $0.01$
   - Norm. BN BN BN BN BN
Graph pooling sum mean mean -- --
PE dim (RW-steps) 21 18 18 21 32
Kernel Func.
   - # MLP Block 2 2 2 2 2
   - Norm. BN BN BN BN BN
   - Kernel dropout $0.5$ $0.5$ $0.5$ $0.5$ $0.5$
   - MLP dropout $0.1$ $0.2$ $0$ $0.2$ $0.5$
Batch size 32 16 16 16 16
Learning Rate $0.001$ $0.001$ $0.001$ $0.001$ $0.001$
# Epochs 2000 200 200 200 200
# Warmup epochs 50 5 5 10 10
Weight decay $51\mathrm{e}-$ 5 $51\mathrm{e}-$ 5 $51\mathrm{e}-$ 5 $51\mathrm{e}-$ 5 $51\mathrm{e}-$ 5
Min. lr. $61\mathrm{e}-$ 6 $41\mathrm{e}-$ 4 $41\mathrm{e}-$ 4 $41\mathrm{e}-$ 4 $41\mathrm{e}-$ 4
# Parameters 433,663 102,580 105,320 438,143 499,754
Table 12: Hyperparameters for two datasets from the Long-range Graph Benchmark Dwivedi et al. [2022b]
Hyperparameter Peptides-func Peptides-struct
# CKGConv-Block 4 4
   - Hidden dim 96 96
   - Dropout 0 0.05
   - Norm. BN BN
Graph pooling mean mean
PE dim (walk-step) 24 24
Kernel Func.
   - # MLP Block 2 2
   - Norm. BN BN
   - Kernel dropout 0.5 0.2
   - MLP dropout 0.2 0.2
Batch size 16 16
Learning Rate 0.001 0.001
# Epochs 200 200
# Warmup epochs 5 5
Weight decay 0 0
Min. lr. 1e-4 1e-4
# Parameters 421,468 412,253
B.5 Significance Test
We conduct a two-sample one-tailed t-test to verify the significance of the performance difference. The baselines’ results are taken from Ma et al. [2023].
The statistical tests are conducted using the tools available at https://www.statskingdom.com/140MeanT2eq.html.
B.6 Runtime
We provide the runtime and GPU memory consumption of CKGCN in comparison to GRIT on ZINC as reference (Table 13).
Table 13: Runtime and GPU memory for GRIT Ma et al. [2023] and CKGCN (Ours) on ZINC with batch size $32$. The timing is conducted on a single NVIDIA V100 GPU (Cuda 11.8) and 20 threads of Intel(R) Xeon(R) Gold 6140 CPU @ 2.30GHz.
ZINC CKGConv GRIT
GPU Memory 2146 MB 1896 MB
Training time 35.9 sec/epoch 39.7 sec/epoch
B.7 Efficient Implementation
A more efficient implementation is achievable for CKGConv with global support when the graph order is large. Based on the following derivation, we can implement an algorithm with $O(|\mathcal{V}|S)$ complexity.
Let $\mathcal{S}_{i}:=\{j\in\mathcal{V}:\mathbf{P}_{i,j}\neq\mathbf{0}\}$, then ignoring the bias term, Eq. (3) can be written as
$(\chi\star\psi)(i)$ $=\frac{1}{|\mathcal{V}|}\left(\sum_{j\in\mathcal{S}_{i}}\chi(j)|\mathcal{V}|\cdot\psi(\mathbf{P}_{i,j})+\sum_{j^{\prime}\in\mathcal{V}\backslash\mathcal{S}_{i}}\chi(j)\cdot\psi(\mathbf{0})\right)\,,$ (11)
$=\frac{1}{|\mathcal{V}|}\left(\sum_{j\in\mathcal{S}_{i}}\chi(j)|\mathcal{V}|\cdot\big{(}\psi(\mathbf{P}_{i,j})-\psi(\mathbf{0})\big{)}\right)+\frac{1}{|\mathcal{V}|}\left(\sum_{j\in\mathcal{V}}\chi(j)\cdot\psi(\mathbf{0})\right)\,,$ (12)
$=\frac{1}{|\mathcal{V}|}\left(\sum_{j\in\mathcal{S}_{i}}\chi(j)|\mathcal{V}|\cdot\big{(}\psi(\mathbf{P}_{i,j})-\psi(\mathbf{0})\big{)}\right)+\psi(\mathbf{0})\cdot\frac{1}{|\mathcal{V}|}\left(\sum_{j\in\mathcal{V}}\chi(j)\right)\,.$ (13)
The second term of Eq. (13) can be computed by global-average pooling of graphs shared by all nodes in $O(|\mathcal{V}|)$, and the first term requires $O(|\mathcal{V}|\cdot S)$ computation on average, where $S=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}|\mathcal{S}_{i}|$.
Appendix C Other Related Work
C.1 Graph Positional Encoding
In recent years, positional and/or structural encoding has been widely studied to enhance the performance of both MPNNs You et al. [2019], Ma et al. [2021], Li et al. [2020], Zhang et al. [2021], Loukas [2020], Dwivedi et al. [2021], Lim et al. [2023], Wang et al. [2022], You et al. [2021], Velingker et al. [2023], Bouritsas et al. [2022] and graph transformers Dwivedi and Bresson [2021], Kreuzer et al. [2021], Ying et al. [2021], Zhang et al. [2023], Ma et al. [2023]. Positional encodings prioritize distance/affinity measures and structural encodings focus on structural patterns, but most encodings incorporate both positional and structural information Srinivasan and Ribeiro [2020].
Appendix D Theory and Proof
D.1 The Expressiveness of CKGConv Is Equivalent to GD-WL
We use a Weisfeiler-Lehman (WL)-like graph isomorphism framework to analyze theoretical expressiveness. Specifically, we consider the Generalized Distance WL (GD-WL) test, which is based on updating node colors incorporating graph distances proposed by Zhang et al. [2023].
For a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, the iterative node color update in GD-WL test is defined as:
$\chi^{\ell}_{\mathcal{G}}(v)=\textit{hash}(\{\!\!\{(d_{\mathcal{G}}(v,u),\chi^{\ell-1}_{\mathcal{G}}(u)):u\in\mathcal{V}\}\!\!\})\,.$ (14)
where $d_{\mathcal{G}}(v,u)$ denotes a distance between nodes $v$ and $u$, and $\chi_{G}^{0}(v)$ is the initial color of $v$. The multiset of final node colors $\{\!\!\{\chi_{G}^{L}(v):v\in\mathcal{V}\}\!\!\}$ at iteration $L$ is hashed to obtain a graph color.
Our proof for the expressiveness of CKGConv employs the following lemma provided by Xu et al. [2019].
Lemma D.1.
(Lemma 5 of  Xu et al. [2019]) For any countable set $\mathcal{X}$, there exists a function $f:\mathcal{X}\to\mathbb{R}^{n}$ such that $h(\hat{\mathcal{X}}):=\sum_{x\in\hat{\mathcal{X}}}f(x)$ is unique for each multiset $\hat{\mathcal{X}}\in\mathcal{X}$ of bounded size. Moreover, for some function $\phi$, any multiset function $g$ can be decomposed as $g(\hat{\mathcal{X}})=\phi(\sum_{x\in\hat{\mathcal{X}}}f(x))$.
Proof of Proposition 3.1.
In this proof, we consider shortest-path distance (SPD) as an example of generalized distance (GD). This is denoted as $\text{d}_{G}^{\text{SPD}}$ and is assumed to construct the pseudo-coordinates in CKGConv. The proof holds with other GDs such as the resistance distance (RD) Zhang et al. [2023] and RRWP Ma et al. [2023], and the choice of GD determines the practical expressiveness of GD-WL.
We consider all graphs with at most $n$ nodes to distinguish in the isomorphism tests. The total number of possible values of $\text{d}_{G}$ is finite and depends on $n$ (upper bounded by $n^{2}$). We define
$\mathcal{D}_{n}=\{\text{d}_{G}^{\text{SPD}}(u,v):G=(\mathcal{V},\mathcal{E}),|\mathcal{V}|\leqslant n,u,v\in\mathcal{V}\}\,,$ (15)
to denote all possible values of $\text{d}_{G}^{\text{SPD}}(u,v)$ for any graphs with at most $n$ nodes. We note that since $\mathcal{D}_{n}$ is a finite set, its elements can be listed as $\mathcal{D}_{n}=\{d_{G,1},\cdots,d_{G,|\mathcal{D}_{n}|}\}$.
Then the GD-WL aggregation at the $\ell$-th iteration in Eq. (14) can be equivalently rewritten as (See Theorem E.3 in Zhang et al. [2023]):
$\chi^{\ell}_{G}(v):=\text{hash}\Big{(}\chi_{G}^{\ell,1}(v),\chi_{G}^{\ell,2}(v),\cdots,\chi_{G}^{\ell,|\mathcal{D}_{n}|}(v)\Big{)}\,,$
where $\chi_{G}^{\ell,k}(v):=\{\!\!\{\chi_{G}^{\ell-1}(u):u\in\mathcal{V},\text{d}_{G}(u,v)=d_{G,k}\}\!\!\}\,.$ (16)
In other words, for each node $v$, we can perform a color update by hashing a tuple of color multisets. We construct the $k$-th multiset by injectively aggregating the colors of all nodes $u\in\mathcal{V}$ at a distance $d_{G,k}$ from node $v$.
Assuming the color of each node $\chi^{t}_{G}(v)$ is represented as a vector $\mathbf{x}_{v}^{(l)}\in\mathbb{R}^{C}$, and setting the bias $\mathbf{b}$ to $\mathbf{0}$ for simplicity, the $l$-th CK-GConv layer with a global support (as shown in Eq. (4)) can be written as
$\hat{\mathbf{x}}^{(l)}_{v}:=\frac{1}{|\mathcal{V}|}\sum_{u\in\mathcal{V}}(\mathbf{W}\mathbf{x}^{(l)}_{u})\odot\boldsymbol{\psi}\big{(}\text{d}_{G}(u,v)\big{)}\,.$ (17)
where $\boldsymbol{\psi}:\mathbb{R}\to\mathbb{R}^{C}$ and $\mathbf{W}\in\mathbb{R}^{C\times C}$ is the learnable weight. Then, we will show that with certain choices of the kernel function, a CKGCN is as powerful as GD-WL.
[…]
Then Eq. (17) can be written as
$\hat{\mathbf{x}}^{(l),h}_{v}:=$ $\frac{1}{|\mathcal{V}|}\sum_{u\in\mathcal{V}}(\mathbf{W}^{h}\mathbf{x}^{(l)}_{u})\odot\boldsymbol{\psi}^{h}\big{(}\text{d}_{G}(u,v)\big{)}\,,$ (18)
$\hat{\mathbf{x}}^{(l)}_{v}=$ $[\hat{\mathbf{x}}^{(l),1}_{v}\|\cdots\|\hat{\mathbf{x}}^{(l),H}_{v}]\,,$ (19)
where $\mathbf{W}\in\mathbb{R}^{C\times C}$ is partitioned as $[{\mathbf{W}^{1}}^{\intercal},\cdots,{\mathbf{W}^{H}}^{\intercal}]^{\intercal}$ so that each $\mathbf{W}^{h}\in\mathbb{R}^{F\times C}$.
We construct $\boldsymbol{\psi}^{h}(d):=\mathbb{I}(d=d_{G,h})\cdot\mathbf{1}$, where $\mathbb{I}:\mathbb{R}\to\mathbb{R}$ is the indicator function, $d_{G,h}\in\mathcal{D}_{n}$ is a pre-determined condition, and $\mathbf{1}\in\mathbb{R}^{F}$. Then, the convolution by each sub-kernel (Eq. (18) can be written as
$\hat{\mathbf{x}}^{(l),h}_{v}:=$ $\frac{1}{|\mathcal{V}|}\sum_{u\in\mathcal{V}}(\mathbf{W}^{h}\mathbf{x}^{(l)}_{u})\odot\boldsymbol{\psi}^{h}\big{(}\text{d}_{G}(u,v)\big{)}\,,$ (20)
$=$ $\frac{1}{|\mathcal{V}|}\sum_{u\in\mathcal{V}}(\mathbf{W}^{h}\mathbf{x}^{(l)}_{u})\odot(\mathbb{I}(\text{d}_{G}(u,v)=d_{G,h})\cdot\mathbf{1}\big{)}\,,$
$=$ $\frac{1}{|\mathcal{V}|}\sum_{u\in\mathcal{V}}(\mathbf{W}^{h}\mathbf{x}^{(l)}_{u})\cdot\mathbb{I}(\text{d}_{G}(u,v)=d_{G,h})\,,$
$=$ $\frac{1}{|\mathcal{V}|}\sum_{\text{d}_{G}(u,v)=d_{G,h}}{\mathbf{W}}^{h}\mathbf{x}^{(l)}_{u}\,.$
Note that ${\mathbf{W}}$ can be absorbed as the last layer of the feed-forward network (FFN) in the previous layer. Because $\mathbf{x}^{(l)}_{u}$ is processed by the FFN in the previous layer, we can invoke Lemma D.1 to establish that each sub-kernel $\boldsymbol{\psi}^{h}$ (as in Eq. (20)) can implement an injective aggregating function for $\{\!\!\{\chi_{G}^{t-1}(u):u\in\mathcal{V},d_{G}(u,v)=d_{G,h}\}\!\!\}$. The concatenation in Eq. (19) is an injective mapping of the tuple of multisets $\left(\chi_{G}^{t,1},\cdots,\chi_{G}^{t,|\mathcal{D}_{n}|}\right)$. When any of the linear mappings has irrational weights, the projection will also be injective. Therefore, one CKGConv followed by the FFN can implement the aggregation formula (Eq. (16)), with a sufficiently large number of different $\boldsymbol{\psi}^{h}$. Thus, the CKGCN can perform the aggregation of GD-WL. Therefore, with a sufficiently large number of layers, CKGCN is as powerful as GD-WL in distinguishing non-isomorphic graphs, which concludes the proof. ∎
D.2 CKGConv and Equivariant Set Neural Networks
Proof of Proposition 4.1.
We prove the proposition for scalar-valued signals, which can be directly generalized to vector-valued signals.
For a globally supported CKGConv, the $1$-order RRWP after the re-normalization of relative positions is $P_{i,i}=|\mathcal{V}|\text{ and }P_{i,j}=0,\forall i,j\in\mathcal{V},i\neq j$, denoted as $P_{0}$ and $P_{1}$ for simplicity. Considering $\psi:\mathbb{R}\to\mathbb{R}$ that $\psi(x)=\gamma\cdot x+\beta$ $\gamma,\beta\in\mathbb{R}$, Eq. (3) can be written as
$(\chi\star\psi)(i)$ $=\frac{1}{|\mathcal{V}|}\left(\chi(i)(|\mathcal{V}|\cdot\gamma+\beta)+\sum_{j\in\mathcal{V};j\neq i}\chi(j)\beta\right)+b\,,$ (21)
$=\frac{1}{|\mathcal{V}|}\chi(i)(|\mathcal{V}|\cdot\gamma+\beta-\beta)+\frac{1}{|\mathcal{V}|}\sum_{j\in\mathcal{V}}\chi(j)\beta+b\,,$
$=\gamma\cdot\chi(i)+\beta\cdot\big{(}\frac{1}{|\mathcal{V}|}\sum_{j\in\mathcal{V}}\chi(j)\big{)}+b\,.$
This is the general form of a layer in an Equivariant Set Network (Eq. 8 in Segol and Lipman [2020]). This general form can cover a wide range of set neural networks Zaheer et al. [2017], Qi et al. [2017]. ∎
D.3 CKGConv, Polynomial Spectral GNNs and Diffusion Enhaned GNNs
Lemma D.2.
Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ denotes the adjacency matrix of an undirected graph $G$ and the diagonal matrix $\mathbf{D}\in\mathbb{R}^{n\times n},[\mathbf{D}]_{i,i}=\sum_{j\in\mathcal{V}}[\mathbf{A}]_{i,j}$ is the degree matrix, the $k$-power of symmetric normalized adjacency matrix $\tilde{\mathbf{A}}:=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ and random walk matrix $\mathbf{M}:=\mathbf{D}^{-1}\mathbf{A}$, satisfy that
$\tilde{\mathbf{A}}^{k}=\mathbf{D}^{1/2}\mathbf{M}\mathbf{D}^{-1/2},\forall k=1,2,\cdots$ (22)
Proof of Lemma D.2.
For arbitrary $k\geq 1$, we have
$\tilde{\mathbf{A}}^{k}$ $=(\mathbf{D}^{1/2}\mathbf{M}\mathbf{D}^{-1/2})^{k}$ (23)
$=(\mathbf{D}^{1/2}\mathbf{M}\mathbf{D}^{-1/2})(\mathbf{D}^{1/2}\mathbf{M}\mathbf{D}^{-1/2})\cdots(\mathbf{D}^{1/2}\mathbf{M}\mathbf{D}^{-1/2})\text{ ,$k$ times}$
$=\mathbf{D}^{1/2}\mathbf{M}\cdots\mathbf{M}\mathbf{D}^{-1/2}$
$=\mathbf{D}^{1/2}\mathbf{M}^{k}\mathbf{D}^{-1/2}\,.$
Proof of Proposition 4.2.
Irrespective of the specific polynomial parameterization that is employed, any $K{-}1$-th order Polynomial Spectral Graph Neural Network can be defined in a general form with $\tilde{\mathbf{L}}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}\in\mathbb{R}^{n\times n}$, parameterized by a learnable vector $\mathbf{\theta}=[\theta_{0},\cdots,\theta_{K-1}]^{\intercal}\in\mathbb{R}^{K\times 1}$ for the filtering of an input graph signal $\mathbf{x}\in\mathbb{R}^{n\times 1}$ to obtain an output graph signal $\mathbf{y}\in\mathbb{R}^{n\times 1}$, as follows:
$\mathbf{y}$ $=g_{\mathbf{\theta}}(\tilde{\mathbf{L}})\mathbf{x}\,,$
$=\sum_{k=0}^{K-1}\theta_{k}\tilde{\mathbf{L}}^{k}\mathbf{x}\,,$
$=\sum_{k=0}^{K-1}\theta_{k}\sum_{r=0}^{k}{k\choose r}(-1)^{r}\tilde{\mathbf{A}}^{r}\mathbf{x}\,,$
$=\sum_{k=0}^{K-1}\theta^{\prime}_{k}\tilde{\mathbf{A}}^{k}\mathbf{x}\,.$ (24)
Here $\tilde{\mathbf{A}}=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ and $\theta^{\prime}_{k}=\sum_{r=k}^{K-1}{k\choose r}(-1)^{r}\theta_{r}$. Therefore, the spectral filter $g_{\mathbf{\theta}}(\tilde{\mathbf{L}})$ can be represented by a linear combination of a collection of polynomial bases $\{\mathbf{I},\tilde{\mathbf{A}}^{1},\tilde{\mathbf{A}}^{2},\cdots,\tilde{\mathbf{A}}^{K-1}\}$. In other words,
i,j $=\psi([\mathbf{I},\tilde{\mathbf{A}}^{1},\tilde{\mathbf{A}}^{2},\cdots,\tilde{\mathbf{A}}^{K-1}]_{i,j})\,,$ (25)
[…]
$=d_{i}^{1/2}\psi([\mathbf{I},\mathbf{M}^{1},\mathbf{M}^{2},\cdots,\mathbf{M}^{K-1}]_{i,j})d_{j}^{-1/2}\,,\text{ as $\psi$ is a linear projection}$
$=d_{i}^{1/2}\psi(\mathbf{P}_{i,j})d_{j}^{-1/2}\,.$
$=\frac{1}{S}\cdot d_{i}^{1/2}\psi(S\cdot\mathbf{P}_{i,j})d_{j}^{-1/2}\,.$
Here $\psi:\mathbb{R}^{K}\to\mathbb{R}$ is a linear projection; $d_{i}=\mathbf{D}_{i,i}\in\mathbb{R}$ is the degree of node $i$; and $S\in\mathbb{R}$ is the scaling term in the support-size invariant convolution design and the RRWP preprocessing.
In other words, with the same order $K{-}1$ of RRWP as pseudo-coordinates, CKGConv with a linear kernel $\psi$ can recover most polynomial spectral GNNs in the form of Eq. (24) irrespective of the specific polynomial parameterization that is used, if $d_{i}^{1/2}$ and $d_{j}^{-1/2}$ are injected properly, for all $i,j\in\mathcal{V}$. The result trivially holds for other Laplacian normalizations (e.g., row-normalized, max-eigenvalue normalized), where different constant multipliers are injected via adaptive degree scalers to $\mathbf{P}_{i,j}$.
Similarly, Polynomial Diffusion Enhanced Graph Neural Networks employing polynomials of $\tilde{\mathbf{A}}$ or its variants also can be represented by Eq. (24). Hence, the result follows. ∎
D.4 Degree Information and Normalization Layers
Normalization layers are essential for deep neural networks. Ma et al. [2023] provide a thorough discussion on the impact of normalization layers on the explicit injected degree information via sum-aggregation or degree scalers, which motivates our choice of BatchNorm Ioffe and Szegedy [2015] over LayerNorm Ba et al. [2016].
Proposition D.3.
Ma et al. [2023] Sum-aggregated node representations, degree-scaled node representations, and mean-aggregated node representations all have the same value after application of a LayerNorm on node representations.
Proof of Proposition D.3.
Regardless the linear transformation in MPNN shared by nodes, we can write the output representation for a node $i$ from a sum-aggregator as $\mathbf{x}^{\text{sum}}_{i}=d_{i}\cdot\mathbf{x}^{\text{mean}}_{i}$, where $d_{i}\in\mathbb{R}$ is the degree of node $i$ and $\mathbf{x}^{\text{mean}}_{i}=[x_{i1},\dots x_{iF}]^{\top}\in\mathbb{R}^{F}$ is the node representation from a mean-aggregator. The layer normalization statistics for a node $i$ over all hidden units are computed as follows:
$\mu^{\text{sum}}_{i}=\frac{1}{F}\sum_{j=1}^{F}x^{\text{sum}}_{ij}=\frac{1}{F}\sum_{j=1}^{F}d_{i}\cdot x^{\text{mean}}_{ij}=\frac{d_{i}}{F}\sum_{j=1}^{F}x^{\text{mean}}_{ij}=d_{i}\cdot\mu^{\text{mean}}_{i}$ (26)
$\sigma^{\text{sum}}_{i}=\sqrt{\frac{1}{F}\sum_{j=1}^{F}(x^{\text{sum}}_{ij}-\mu^{\text{sum}})^{2}}=\sqrt{\frac{d_{i}^{2}}{F}\sum_{j=1}^{F}(x^{\text{mean}}_{ij}-\mu^{\text{mean}})^{2}}=d_{i}\cdot\sigma_{i}^{\text{mean}}$
Therefore, regardless of the elementwise affine transforms shared by all nodes, each element of the normalized representation
$\tilde{x}_{ij}^{\text{sum}}=\frac{({x}_{ij}^{\text{sum}}-\mu_{i}^{\text{sum}})}{\sigma^{\text{sum}}_{i}}=\frac{(d_{i}\cdot{x}_{ij}^{\text{mean}}-d_{i}\cdot\mu_{i}^{\text{mean}})}{d_{i}\cdot\sigma^{\text{mean}}_{i}}=\frac{({x}_{ij}^{\text{mean}}-\mu_{i}^{\text{mean}})}{\sigma^{\text{mean}}_{i}}=\tilde{x}_{ij}^{\text{mean}},\quad\forall i\in\mathcal{V},\forall j=1,\dots,F,$ (27)
will be the same for both sum-aggregation and mean-aggregation.
The same conclusion can be seen for degree scalers, by simply changing $d_{i}$ to $f(d_{i})$ in the proof, where $f:\mathbb{R}\to\mathbb{R}_{>0}$. ∎
Note that, BatchNorm will not have such an impact on degree information, since the normalization statistics are computed across all nodes (with different degrees) in each mini-batch per channel.